Two spherical objects with a mass of 9.31 kg each are placed at a distance of 3.75 m apart. How many electrons need to leave each object so that the net force between them becomes zero?
Answer #1

As you know the gravitational force is attractive force .

And electric force is repulsive force ( when objects have same nature charge.

For resultant force will be zero

When. Gravitational force =electric force

Gmm/r^2 = KQQ /r^2

Q^2=Gm^2/K. Here G=6.672×10^(-11),K=9×10^9

m=9.31 kg

Q^2=6.672×10^(-11)×(9.31)^2/9×10^9

Q^2=64.26×10^(-20)

Q=±8.02×10^(-10) C

We know that

Q=ne =8.02×10^(-10)

n=8.02×10^(-10)/1.6×10^(-19)

n=5.0125×10^9 electrons . Answer.